Solve using any method. (x/2) + (y/3) = 2/3 and (x/5) + (5y/2) = 1/4
I cannot figure out how to do this problem I have been working on it for an hour and have not figured it out please help.

Answer by ewatrrr(24785)

 
Hi
Solve using any method.
(x/2) + (y/3) = 2/3 |multiplying thru by 6
(x/5) + (5y/2) = 1/4 |multiplying thru by 20
(3x) + (2y) = 4
(4x) + (50y) = 5
(-4x) + (-8/3y) = 4*(-4/3) = -16/3 |Multiplying 1st EQ by -4/3
(4x) + (50y) = 5
142/3 y= -1/3
y = -1/142 and x = 190/142